OpenStudy (anonymous):

solve -4x-6y=68, x+2y=-21by using substitution. Express your answer as an ordered pair

OpenStudy (anonymous):

Solve one of the equations for x or y. Plug it back into the same equation or the other one.

OpenStudy (anonymous):

I always preferred adding the equations. \[4*(eq_2) + eq_1 \rightarrow 2y = 68-84\ ...\ etc\]

OpenStudy (anonymous):

ok so what we can do here is the following: with equation x + 2y = -21 will become x = - 21 - 2y so from here we can substituate x into the first equation. -4x-6y=68 so -4(-21-2y)-6y = 68 84 + 8y - 6y = 68 2y = -16 so y = -8. so in this case x = -21 - 2(-8) x = -21 + 16 = 5 so x = 5 and y = -8 is the solution to the two equations